blog:20241107

20241107

1830 - 0130

Connor, John, Roy, Brayden, Percy, Joyce

401, 433c, Dome

  • Test Variable Stars Observing Plan
Scope Camera Focus
MainASI 2600MM Pro1976.3
FinderASI 482
Time Observed Source Name Filter Int Time [s] Gain Airmass Mag (filter) RA [h] DEC [deg] Source Type Scope Purpose Notes
20241107THHMM HD 213306 V 6 150 1.1 3.49(minV) 22 58 Cepheid Main Variables No Guiding
20241107THHMM HD 203156 V 14 150 airmass 5.77(minV) 21 38 Cepheid Main Variables No Guiding
20241107THHMM HD 14662 V 60 150 airmass 6.18(minV) 2 55 Cepheid Main Variables No Guiding
20241107THHMM HD 198726 V 30 150 airmass 5.41(minV) 20 28 Cepheid Main Variables No Guiding
20241107THHMM HD 2313306 ? 6 150 airmass HH DD Cepheid Main Variables No Guiding
20241107THHMM HD 213306 ? 6 150 airmass 3.49(minV) 22 58 Cepheid Main Variables No Guiding

Evening Shift

  1. Begin startup procedures
  2. Attempted pointing
    1. Pointing was off from the start, attempted astrometry
    2. Pro Tip: Use epoch J2000 for the coordinates given by astrometry
    3. DC being slow and clunky opening finderscope
    4. Restarted Computer to fix
    5. Computer having trouble finding 2600 now
      1. Reseated 2600 - Appeared to work
  3. We in business >:)
  4. Slewed to del Cep (HD 213306)
    1. Found reference star
    2. GOOD FOR VARIABLES STUDY
    3. 6 second exposures
  5. Slewed to DT Cyg (HD 201078)
    1. Found reference star
    2. Reference too dim for variable study
  6. Slewed to eta Aql (HD 187929)
    1. Found reference star
    2. Reference really dim
      1. Source: 165
      2. Noise: 30
  7. Slewed to V1334 Cyg (HD 203156)
    1. Found reference star
    2. Reference star really dim
    3. 14 second exposure
  8. Slewed to V0440 Per (HD 14662)
    1. Found Reference star
    2. Really dim (but possible?)
    3. 60-second exposure
  9. Slewed to T Vul (HD 198726)
    1. Found Reference Star
    2. DIM:
      1. Reference: 133
      2. Noise 50
    3. 30-second exposure
  10. Slewed to del Cep again
    1. took more pics

Midnight Shift

  1. Took captures of HD 213306
    1. 2300 6s exposure
    2. 0000 6s exposure
  1. shut down start at 0110
  2. completed by 0126

0050 noticed that the filter wheel was unconnected and most likely not resting on V, data most likely not usable so no flats or darks taken

Captures left in local fiels so that Connor can do w them as he likes